OpenStudy (anonymous):

slove the linear equestion 2x+3y=1, 5X+7y=3. 2) 2X+4y=10 and 3x+6y=15 3) 2x+3y=3, x-2=5 and 3x+2y=7

OpenStudy (texaschic101):

2x + 3y = 1 -->(-5) 5x + 7y = 3 -->(2) ---------------- -10x - 15y = -5 (result of multiplying by -5) 10x + 14y = 6 (result of multiplying by 2) ----------------add -y = 1 y = -1 now sub -1 in for y in either equation to find x. ====================== 2x + 4y = 10 reduces to x + 2y = 5 3x + 6y = 15 reduces to x + 2y = 5 same equation, same line, INFINITE SOLUTIONS ====================== In these 3 equations....is there supposed to be a y term in the 2nd one ?

OpenStudy (anonymous):

yep